Aubree:

When the movie is half over, popcorn sells for half price. Full price popcorn is $3.50, and drinks are $1.35. Determine the cost for purchasing 3 orders of popcorn and 3 drinks when the movie is half over.

TWALKER16:

1.75*3=$5.25, 1.35*3=$4.05, so 5.25+4.05=?

Aubree:

This is the expression from the last question

Aubree:

1 attachment
10kbands:

um ok

Aubree:

When I do that, I get 9.3

10kbands:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Aubree When I do that, I get 9.3 \(\color{#0cbb34}{\text{End of Quote}}\) ummm i don't know lol

Aubree:

@mxddi3

mxddi3:

ok so post the image with the answer choices as well, please

Aubree:

No answer choses

Aubree:

1 attachment
mxddi3:

no, show me the image you sent me in my dms....

Aubree:

That was a different one, brother helped

Aubree:

For that one, it was the top one

mxddi3:

yes, i know ok anyways you know the equation you had for that other question? the: \[3[\frac{ 1 }{ 2 }(3.50)+1.35)]\]

Aubree:

Yes

mxddi3:

this is literally the same question

Aubree:

No

mxddi3:

this is what you sent me....it's asking if popcorn is at half price, what is the cost of 3 popcorns and 3 drinks....it's the same thing

mxddi3:

which is literally the same question you asked at the top

Aubree:

This one I need to find the answer. The last one I need the expression for this one to salve it

Aubree:

Solve

mxddi3:

but i'll explain it to you, again. since the popcorn is selling for half price and it was originally $3.50, then you need to do 1/2(3.50) and drinks are 1.35 so you need to add these together. 1/2(3.50)+1.35 but, since you are buying 3 of each, you need to multiply the entire thing by 3. so you literally do 1/2(3.50)=1.75 and then you have 3(1.75+1.35) and i'm sure you can do that yourself, as i already told you....

mxddi3:

^use the same equation from earlier

Aubree:

This is what I got

1 attachment
mxddi3:

is there anything after the 9? there's a glare lmaooo

Aubree:

It's 9.3

mxddi3:

yes

Aubree:

So that but add a 0

mxddi3:

yuhh